# Ancora: Accurate Intrusion Recovery for Web Applications

### v2

@@ VI Related Work
  • Taint-based Analysis. Systems like Shuttle [16] use deep instrumentation of the application runtime to propagate taint markers from http requests to low-level operations. This approach, while precise in theory, is highly invasive, making it difficult to deploy and maintain in complex production applications. It also fails to track operations from unexpected execution paths created by vulnerabilities like RCE. In contrast, Ancora uses non-invasive, framework-level instrumentation and syscall-based causal analysis to cover all execution paths.
  • Black-box Correlation. Systems like Sanare [11] and Rectify [17] treat the application as a black box, using machine learning to correlate high-level requests with low-level system events. Their accuracy degrades significantly under high concurrency due to the noisy and interleaved nature of system events. Furthermore, their file monitoring is often restricted to predefined directories. Ancora achieves superior accuracy through deterministic partitioning and attribution, and its provenance analysis provides whole-system file tracking.
  • System-wide Causal Tracking. Approaches like Retro [39] build a global graph of system actions to support recovery. In a high-concurrency web environment, this leads to dependency explosion[23, 25, 26, 24, 22, 27], creating a massive, tangled graph that is intractable to analyze. Ancora’s core design overcomes this by first partitioning the global log into independent, request-specific causal chains, then using a targeted method to solve the most complex dependency (the database), thus avoiding the need to analyze a monolithic graph.
  • Database-specific Recovery. Work such as Warp [14] focuses on fine-grained database recovery by rolling back and selectively replaying non-malicious transactions. However, these systems assume that the malicious transactions are already identified and do not address the challenge of attributing transactions to a specific web request in modern architectures with connection pooling. Ancora’s key contribution is solving this attribution problem with our spatiotemporal anchor method, providing the missing link that makes rollback-and-replay strategies like Warp’s practical in real-world settings.

### v3

@@ VI Related Work
  • Taint-based Analysis. Systems like Shuttle [nascimento2015shuttle] use deep instrumentation of the application runtime to propagate taint markers from http requests to low-level operations. This approach, while precise in theory, is highly invasive, making it difficult to deploy and maintain in complex production applications. It also fails to track operations from unexpected execution paths created by vulnerabilities like RCE. In contrast, Ancora uses non-invasive, framework-level instrumentation and syscall-based causal analysis to cover all execution paths.
  • Black-box Correlation. Systems like Sanare [matos2021sanare] and Rectify [matos2017rectify] treat the application as a black box, using machine learning to correlate high-level requests with low-level system events. Their accuracy degrades significantly under high concurrency due to the noisy and interleaved nature of system events. Furthermore, their file monitoring is often restricted to predefined directories. Ancora achieves superior accuracy through deterministic partitioning and attribution, and its provenance analysis provides whole-system file tracking.
  • System-wide Causal Tracking. Approaches like Retro [kim2010intrusion] build a global graph of system actions to support recovery. In a high-concurrency web environment, this leads to dependency explosion[lee2013high, ma2017mpi, alhanahnah2022autompi, yu2021alchemist, wang2023tesec, hassan2020omegalog], creating a massive, tangled graph that is intractable to analyze. Ancora’s core design overcomes this by first partitioning the global log into independent, request-specific causal chains, then using a targeted method to solve the most complex dependency (the database), thus avoiding the need to analyze a monolithic graph.
  • Database-specific Recovery. Work such as Warp [chandra2011intrusion] focuses on fine-grained database recovery by rolling back and selectively replaying non-malicious transactions. However, these systems assume that the malicious transactions are already identified and do not address the challenge of attributing transactions to a specific web request in modern architectures with connection pooling. Ancora’s key contribution is solving this attribution problem with our spatiotemporal anchor method, providing the missing link that makes rollback-and-replay strategies like Warp’s practical in real-world settings.